There are 850 Douglas fir and ponderosa pine trees in a section of forest bought by Sawz Logging Co. The company paid an average of $300 for each Douglas fir and $225 for each ponderosa pine. If the company paid $217,500 for the trees, how many of each kind did the company buy?
Answer by checkley71(8403)

D+P=850 D=850-P
300D+225P=217,500
NOW REPLACE D WITH(850-P) & SOLVE FOR D
300(850-P)+225P=217,500
255,000-300P+225P=217,500
-75P=217,500-255,000
-75P=-37,500
P=-500/-75
P=500 NUMBER OF PONDEROSA PINES SOLD.
850-500=350 DOUGLAS FIRS.
PROOF
300*350+225*500=217,500
105,000+112,500=217,500
127,500=217,500